OpenStudy (nickymail48):

3x-7y=4 5x-3y=4 solve with addition method

OpenStudy (anonymous):

15*x-35*y=20 15*x-9*y=12 -26*y=8 y=8/(-26) x=0.615

OpenStudy (nickymail48):

i still dont understand

OpenStudy (michele_laino):

Please, if I multiply both sides of the first equation by 5, I get: \[5*(3x-7y)=5*4\] or (1): \[15x-35y=20\] if I multiply both sides of the second equation by -3, I get: \[-3*(5x-3y)=-3*4\] or: (2) \[-15x+9y=-12\] now adding both sides of equation (1), with both sides of equation (2), we get: \[(15x-35y)+(-15x+9y)=20+(-12)\] or:(3) \[-26y=8\] dividing both sides of equation (3) by -26, we get: \[y=\frac{ -26 }{ 8 }=-\frac{ 13 }{ 4 }\]

OpenStudy (michele_laino):

sorry the lastequation are non correct, I rewrite it: \[y=-\frac{ 8 }{ 26 }=-\frac{ 4 }{ 13 }\]

OpenStudy (michele_laino):

next steo is to substitute y=-4/13, for example in the second eqation, namely : 5x-3y=4, so I can write: \[5x-3*(-\frac{ 4 }{ 13 })=4\] or:(4) \[5x+\frac{ 12 }{ 13 }=4\] mltipling both sides of equation (4) by 13, we get: \[65x+12=52\] then: \[65x=40\] dividig both sides of the above equation,by 65, we can write: \[x=\frac{ 40 }{ 65 }=\frac{ 8 }{ 13 }\]

OpenStudy (nickymail48):

thanks

OpenStudy (michele_laino):

thanks!